OpenStudy (anonymous):

Solve the equation for s: -18s + 24 = -9s – 12

OpenStudy (anonymous):

9

OpenStudy (studygurl14):

\(\large\tt\color{blue}{-18s}\color{red}{+24}=\color{blue}{-9s}\color{red}{-12}\) Combine like terms (same-colored numbers) \(\large\tt\color{blue}{-18s+9s}=\color{red}{-12-24}\) \(\large\tt -9s=-36\) \(\large\tt s=4\)

OpenStudy (studygurl14):

And WELCOME TO OPENSTUDY, @susannlovee :)

OpenStudy (anonymous):

@StudyGurl14 thanks :)

OpenStudy (anonymous):

\[Select the ordered pair that is the solution of this equation: -7x – 4y = 6\]

OpenStudy (studygurl14):

You're welcome @susannlovee To show you're appreciation, on OpenStudy, there is a medal system. To give a medal to someone who helped you, click "best response"